OpenStudy (anonymous):

Ruth is having hardwood floors installed in her living room. The cost for the material is $2840 plus an installation charge of $1.90 per square foot. If the total cost for the material plus installation is $5120, determine the area of her living room.

OpenStudy (anonymous):

1200ft^2

OpenStudy (anonymous):

How did you get that

OpenStudy (anonymous):

a= 5120-2840/1.90 = 1200ft^2

OpenStudy (anonymous):

Thanks man

OpenStudy (anonymous):

5120 = 2840 + 1.9x 2280 = 1.9x 1200 = x